# Yibin Wang 1 &Haizhou Shi ∗2 &Ligong Han 2

### v2

@@ 3.4 On Improving the Sample Efficiency of BLoB
15:     Updater the parameters:  ${\bm{M}}\leftarrow{\bm{M}}-\eta\Delta_{{\bm{M}}}$;  ${\bm{G}}\leftarrow{\bm{G}}-\eta\Delta_{{\bm{G}}}$;  ${\bm{B}}\leftarrow{\bm{B}}-\eta\Delta_{{\bm{B}}}$
@@ 4.1 Settings
Baselines and Implementation Details. We compare BLoB with state-of-the-art uncertainty estimation methods applied to the LoRA adapters of LLMs, including Maximum Likelihood Estimation (MLE[41], Maximize A Posteriori (MAP), Monte-Carlo Dropout (MCD[29], Bayes By Backprop (BBB[11], Deep Ensemble (ENS[51, 8, 96], and the latest Laplace-LoRA (LAP[108].
For MLE, we use the LoRA implementation. For MAP, we use a weight decay rate of $51e-$ 5. For MCD, we use an ensemble of 10 LoRAs with a dropout rate of $p=0.1$. For ENS, we independently fine-tune 3 LoRAs and average their logits during evaluation. For BBB, we adopt the default settings from the Bayesian-Torch library [48] and only Bayesianize the ${\bm{A}}$ matrix, similar to BLoB. We re-implement LAP and apply it to the MAP checkpoints. We keep all BLoB-specific hyperparameters consistent across all datasets. Typically, we set the number of samples $K=1$ during training for all our BLoB experiments, which highlights BLoB’s sampling efficiency. As shown in Table 1, we also report BLoB’s performance with different numbers of samples during Bayesian inference, where $N=0$ indicates directly using the mean of the weight distribution for prediction.
@@ A.1 Proof of Main Theorems
It is crucial to note here, the final covaraince matrix of $q(\operatorname{vec}({\bm{W}}))$ follows a block-diagonal structure, which will be further utilized for the proof of Theorem 3.2. Defining ${\bm{\Sigma}}_{i}=\operatorname{diag}({\bm{\Omega}}_{i:}^{2})$, we have:
@@
Another important fact about ${\bm{\Sigma}}_{q}$ is its singularity. It can be see directly as we consider the rank of any one of the block matrix ${\bm{B}}{\bm{\Sigma}}_{i}{\bm{B}}^{\top}\in\mathbb{R}^{m\times m},\forall i\in[n]$:
@@
First, we calculate the log-determinant part of the KL-divergence. For the log-determinant of the covaraince matrix of the prior distribution ${\bm{\Sigma}}_{p}$, by applying SVD decomposition in Eqn. 28, we have
@@ C.4 Embedding Uncertainty of BLoB: A Preliminary Visual Study
Estimating the uncertainty of LLMs in the embedding space has recently garnered significant attention in the community [13]. Expressing models’ uncertainty via their generated embeddings can benefit both discriminative (the focus of this paper) and generative models. In this section, we present a preliminary study on uncertainty estimation in the embedding spaces of different models, as illustrated in Fig. 3. We compare BLoB with two baseline models, BBB and MCD, which can generate embedding samples and effectively estimate uncertainty. We exclude LAP from this section due to its excessive memory consumption, which consistently results in Out-Of-Memory (OOM) errors during inference.
For each input sequence ${\bm{s}}$, we use the last token’s embedding generated by the final transformer block in Llama2-7B as the final embedding. Given the weights ${\bm{W}}$, we denote the embedding as $\phi({\bm{s}};{\bm{W}})$. Generally, three types of embeddings can be generated using the Bayesian approach:
@@
    $\phi({\bm{s}};\mathbb{E}_{{\bm{W}}\sim q(\cdot|{\bm{\theta}})}[{{\bm{W}}}])$ $=\phi({\bm{s}};{\bm{W}}_{0}+{\bm{B}}{\bm{M}}).$ (75)
@@
    $\mathbb{E}_{{\bm{W}}\sim q(\cdot|{\bm{\theta}})}[\phi({\bm{s}};{\bm{W}})]$ $\approx\tfrac{1}{N}\sum_{n=1}^{N}\phi({\bm{s}};{\bm{W}}_{0}+{\bm{B}}({\bm{M}}+{\bm{E}}_{n}\circ{\bm{\Omega}})),$ (76)
@@
Figure 3: Visualization of embedding uncertainty quality for different methods. The model is fine-tuned for 5,000 steps on the Llama2-7B. The two contour lines represent the probability mass of 0.5 and 0.75, respectively.

### v3

# Yibin Wang ${}^{\text{\textdagger}1}$ &Haizhou Shi ${}^{*\text{\textdagger}1}$ &Ligong Han 12

@@ 3.4 On Improving the Sample Efficiency of BLoB
15:     Update the parameters:  ${\bm{M}}\leftarrow{\bm{M}}-\eta\Delta_{{\bm{M}}}$;  ${\bm{G}}\leftarrow{\bm{G}}-\eta\Delta_{{\bm{G}}}$;  ${\bm{B}}\leftarrow{\bm{B}}-\eta\Delta_{{\bm{B}}}$
@@ 4.1 Settings
Baselines and Implementation Details. We compare BLoB with state-of-the-art uncertainty estimation methods applied to the LoRA adapters of LLMs, including Maximum Likelihood Estimation (MLE[41], Maximum A Posteriori (MAP), Monte-Carlo Dropout (MCD[29], Bayes By Backprop (BBB[11], Deep Ensemble (ENS[51, 8, 96], and the latest Laplace-LoRA (LAP[108].
For MLE, we use the LoRA implementation. For MAP, we use a weight decay rate of $51e-$ 5. For MCD, we use an ensemble of 10 LoRAs with a dropout rate of $p=0.1$. For ENS, we independently fine-tune 3 LoRAs and average their logits during evaluation. For BBB, we adopt the default settings from the Bayesian-Torch library [48] and only Bayesianize the ${\bm{A}}$ matrix, similar to BLoB. We sample $N=10$ times for BBB during test. We re-implement LAP and apply it to the MAP checkpoints. We keep all BLoB-specific hyperparameters consistent across all datasets. Typically, we set the number of samples $K=1$ during training for all our BLoB experiments, which highlights BLoB’s sampling efficiency. As shown in Table 1, we also report BLoB’s performance with different numbers of samples during Bayesian inference, where $N=0$ indicates directly using the mean of the weight distribution for prediction.
@@ A.1 Proof of Main Theorems
It is crucial to note here, the final covariance matrix of $q(\operatorname{vec}({\bm{W}}))$ follows a block-diagonal structure, which will be further utilized for the proof of Theorem 3.2. Defining ${\bm{\Sigma}}_{i}=\operatorname{diag}({\bm{\Omega}}_{i:}^{2})$, we have:
@@
Another important fact about ${\bm{\Sigma}}_{q}$ is its singularity. It can be seen directly as we consider the rank of any one of the block matrix ${\bm{B}}{\bm{\Sigma}}_{i}{\bm{B}}^{\top}\in\mathbb{R}^{m\times m},\forall i\in[n]$:
@@
First, we calculate the log-determinant part of the KL-divergence. For the log-determinant of the covariance matrix of the prior distribution ${\bm{\Sigma}}_{p}$, by applying SVD decomposition in Eqn. 28, we have
@@ C.4 Embedding Uncertainty of BLoB: A Preliminary Visual Study
Estimating the uncertainty of LLMs in the embedding space has recently garnered significant attention in the community [13]. Expressing models’ uncertainty via their generated embeddings can benefit both discriminative (the focus of this paper) and generative models. In this section, we present a preliminary study on uncertainty estimation in the embedding spaces of different models, as illustrated in Fig. 3. We compare BLoB with two baseline models, BBB and MCD, which can generate embedding samples and effectively estimate uncertainty. We exclude LAP from this section due to its excessive memory consumption, which consistently results in Out-Of-Memory (OOM) errors during inference. The experiment is conducted on the OBQA dataset [64], which consists of four categories.
For each input sequence ${\bm{s}}$, we use the last token’s embedding generated by the final transformer block in Llama2-7B as the final embedding. Given the weights ${\bm{W}}$, we denote the embedding as ${\bm{\phi}}({\bm{s}};{\bm{W}})$. Generally, three types of embeddings can be generated using the Bayesian approach:
@@
    ${\bm{\phi}}({\bm{s}};\mathbb{E}_{{\bm{W}}\sim q(\cdot|{\bm{\theta}})}[{{\bm{W}}}])$ $={\bm{\phi}}({\bm{s}};{\bm{W}}_{0}+{\bm{B}}{\bm{M}}).$ (75)
@@
    $\mathbb{E}_{{\bm{W}}\sim q(\cdot|{\bm{\theta}})}[{\bm{\phi}}({\bm{s}};{\bm{W}})]$ $\approx\tfrac{1}{N}\sum_{n=1}^{N}{\bm{\phi}}({\bm{s}};{\bm{W}}_{0}+{\bm{B}}({\bm{M}}+{\bm{E}}_{n}\circ{\bm{\Omega}})),$ (76)
@@
Figure 3: Visualization of embedding uncertainty quality for different methods. The model is fine-tuned for 5,000 steps on the Llama2-7B. We fine-tune the Llama2-7B model on the OBQA dataset for 5000 steps. The two contour lines represent the probability mass of 0.5 and 0.75, respectively.